### v1

@@ 5.4 Robust to Reward Function
Because the benchmark utilizes a default reward value of Eq. 1, we set it to $\mathbb{1}\left[\phi(s_{i})=g\right]$ to assess the algorithm’s robustness to variations in reward structure. The comparison is shown in Table 3. From Table 3, in the all 5 goal-conditioned RL benchmarks, we can see Q-WSL still attains significantly higher performance than prior baselines and HER principle is also still important to learning sparse reward. Moreover, DDPG+HER is poor performance compared to GCWSL as it is always zero in indicator reward, this means Q-learning is less robust to reward function. This phenomenon and shortcomings of Q-learning have also been demonstrated in (Eysenbach et al., 2020).
@@ 5.5 Ablation Studies
To evaluate the significance of incorporating both Q-learning and advantage-weighted regression in Q-WSL, we conducted ablation experiments comparing various Q-WSL variants with HER. For the joint optimization in Equation Eq. 4.2, the parameter $\eta$ is set to 0.1 by default. We conduct our experiments under the following conditions:
@@
  • No WSL which is equivalent to Q-learning as described in Eq. 8.
Empirical results are shown in Fig. 8.
Figure 8: Ablation studies in FetchPick and FetchPush tasks. The results indicate that Dynamic Programming plays a crucial role in the effectiveness of GCWSL.
As depicted in Fig. 8, Q-learning and GCWSL are essential components. Q-WSL demonstrates faster learning than DDPG+HER, whereas the state-of-the-art baseline DWSL fails to learn in FetchPick and FetchPush, suggesting that supervised learning is sub-optimal for relabeled data. The integration of Q-learning and GCWSL significantly enhances performance.
@@ A.1 PROOF of Theorem 1
$\ \arg\max_{\pi}=\int_{g}\rho_{\mathcal{B}_{r}}(g)\int_{s}d_{\mathcal{B}_{r}}(s|g)\int_{a}\left[A^{\pi}(s,a,g)\right]\boldsymbol{d}a\boldsymbol{d}s\boldsymbol{d}g,$
$\;\int_{a}\pi(a|s,g)\boldsymbol{d}a=1,\;\mathbf{s.t.}~{}\operatorname{KL}(\pi_{\theta}\|\pi_{relabel})\leq\epsilon,$ (12)
where $\rho_{{\mathcal{B}_{r}}}(g)$ represents the goal distribution in relabeled data, and $d_{\mathcal{B}_{r}}(s|g)=\sum_{t=0}^{\inf}\gamma^{t}p(s_{t}=s|\pi_{relabel},g)$ denotes the goal-conditioned un-normalized discounted state distribution induced by the relabeling policy $\pi_{relabel}$ (Sutton and Barto, 2018). In this context, $\pi_{\psi}$ is the learned policy, ${\mathcal{B}_{r}}$ refers to the relabeled data, and $\pi_{relabel}$ signifies the behavioral policy responsible for collecting the dataset. KL represents the KL-divergence measurement metric, and $\epsilon$ is the threshold. By enforcing the Karush-Kuhn-Tucker (KKT) condition on Eq. A.1, we obtain the Lagrangian:
$\ \mathcal{L}(\lambda,\pi,c\alpha)$ (13)
$=\int_{g}\rho_{{\mathcal{B}_{r}}}(g)\int_{s}d_{{\mathcal{B}_{r}}}(s|g)\int_{a}\pi(a|s,g)\left[A^{\pi}(s,a,g)\right]\boldsymbol{d}a\boldsymbol{d}s\boldsymbol{d}g$
@@
$\ Z(s,g)=\exp(-\frac{1}{p_{\mathcal{B}_{r}}(s,g)}\frac{\alpha}{\lambda}-1)$
$=\int_{a}^{\prime}\pi_{\mathcal{B}_{r}}(a^{\prime}|s,g)\exp(\frac{1}{\lambda}A^{\pi}(s,a^{\prime},g))\boldsymbol{d}a^{\prime}.$ (17)
@@
$=\arg\min_{\pi}\mathbb{E}_{s,g\sim{\mathcal{B}_{r}}}\mathbb{E}_{a\sim{\mathcal{B}_{r}}}\big{[}-\pi(a|s,g)\exp(\frac{1}{\lambda}A^{\pi}(s,a,g))\big{]}$ (19)
@@
$\ \mathcal{J}_{wbc}(\pi)=\mathbb{E}_{s,a,g\sim{\mathcal{B}_{r}}}\left[\exp(\frac{1}{\lambda}A^{\pi}(s,a,g)\cdot\log(\pi(a|s,g)))\right]$ (21)
@@ A.3 Baseline Details
    $\ \min_{\pi}-\mathbb{E}_{(s_{t},a_{t},g)\thicksim\mathcal{B}_{r}}\big{[}\gamma^{i-t}\exp_{clip}(A(s_{t},a_{t},g))$
    $\cdot\epsilon(A(s_{t},a_{t},g))\log\pi(a_{t}\mid s_{t},g)\big{]},$ (24)
@@
    $\ \hat{d}(s_{i+1},g)=-\alpha\log\mathbb{E}_{k}\sim p_{\zeta}^{r}(\cdot|(s_{i+1},g)$
    $\cdot\left[e^{-k/(I\alpha)}\right],$ (30)
@@
    where $c(s_{i},g)=1\left[\phi(c_{i+1})\neq g\right]/I$.
@@
    , where $\omega$ is the temperature parameter, and $\pi$ is the policy parameterized by $\theta$.
@@ A.4 Additional Experiments Results
Relabeling Ratio The multi-goal framework we consider assumes that data is associated with goal labels. We explore the influence of the relabeling ratio on performance, as depicted in Fig. 9. The results in Fig. 9 indicate that Q-WSL remains robust across different relabeling ratios. Furthermore, Q-WSL consistently outperforms competitive methods such as WGCSL and DDPG+HER at various relabeling ratios.
@@
The Impact of Hyperparameter $\eta$ Since our approach optimizes both supervised learning (SL) and reinforcement learning (RL) concurrently, this section explores the influence of the balancing parameter $\eta$. We evaluate $\eta$ values from the set $\{0.1,0.2,1.0,3.0\}$ and compare the results against competitive HER-based algorithms such as WGCSL and DDPG+HER, as shown in Fig. 10. The findings in Fig. 10 reveal that Q-WSL consistently delivers superior performance over the other algorithms, regardless of the $\eta$ parameter variation. This demonstrates that our method maintains robustness and is not significantly affected by changes in the $\eta$ parameter.

### v2

@@ 5.4 Robust to Reward Function
Because the benchmark utilizes a default reward value of Equation 1, we set it to $\mathbb{1}\left[\phi(s_{i})=g\right]$ to assess the algorithm’s robustness to variations in reward structure. The comparison is shown in Table 3. From Table 3, in the all 5 goal-conditioned RL benchmarks, we can see Q-WSL still attains significantly higher performance than prior baselines and HER principle is also still important to learning sparse reward. Moreover, DDPG+HER is poor performance compared to GCWSL as it is always zero in indicator reward, this means Q-learning is less robust to reward function. This phenomenon and shortcomings of Q-learning have also been demonstrated in (Eysenbach et al., 2020).
@@ 5.5 Ablation Studies
To evaluate the significance of incorporating both Q-learning and advantage-weighted regression in Q-WSL, we conducted ablation experiments comparing various Q-WSL variants with HER. For the joint optimization in Equation 8, the parameter $\eta$ is set to 0.1 by default.
Figure 5: Ablation studies in FetchPick and FetchPush tasks.
We conduct our experiments under the following conditions:
@@
  • No WSL which is equivalent to Q-learning as described in Equation 6.
Empirical results are shown in Figure 5. As depicted in Figure 5, Q-learning and GCWSL are essential components. Q-WSL demonstrates faster learning than DDPG+HER, whereas the state-of-the-art baseline DWSL fails to learn in FetchPick and FetchPush, suggesting that supervised learning is sub-optimal for relabeled data. The integration of Q-learning and GCWSL significantly enhances performance.
@@ A.1 PROOF of Theorem 1
$\arg\max_{\pi}=\int_{g}\rho_{\mathcal{B}_{r}}(g)\int_{s}d_{\mathcal{B}_{r}}(s|g)\int_{a}\left[A^{\pi}(s,a,g)\right]\boldsymbol{d}a\boldsymbol{d}s\boldsymbol{d}g,$ $\int_{a}\pi(a|s,g)\boldsymbol{d}a=1,$
$\mathbf{s.t.}~{}\operatorname{KL}(\pi_{\theta}\|\pi_{relabel})\leq\epsilon,$ (10)
where $\rho_{{\mathcal{B}_{r}}}(g)$ represents the goal distribution in relabeled data, and $d_{\mathcal{B}_{r}}(s|g)=\sum_{t=0}^{\inf}\gamma^{t}p(s_{t}=s|\pi_{relabel},g)$ denotes the goal-conditioned un-normalized discounted state distribution induced by the relabeling policy $\pi_{relabel}$ (Sutton and Barto, 2018). In this context, $\pi_{\psi}$ is the learned policy, ${\mathcal{B}_{r}}$ refers to the relabeled data, and $\pi_{relabel}$ signifies the behavioral policy responsible for collecting the dataset. KL represents the KL-divergence measurement metric, and $\epsilon$ is the threshold. By enforcing the Karush-Kuhn-Tucker (KKT) condition on Section A.1, we obtain the Lagrangian:
$\mathcal{L}(\lambda,\pi,\alpha)$ $=\int_{g}\rho_{{\mathcal{B}_{r}}}(g)\int_{s}d_{{\mathcal{B}_{r}}}(s|g)\int_{a}\pi(a|s,g)\left[A^{\pi}(s,a,g)\right]\boldsymbol{d}a\boldsymbol{d}s\boldsymbol{d}g$
@@
$Z(s,g)=\exp(-\frac{1}{p_{\mathcal{B}_{r}}(s,g)}\frac{\alpha}{\lambda}-1)\\ =\int_{a}^{\prime}\pi_{\mathcal{B}_{r}}(a^{\prime}|s,g)\exp(\frac{1}{\lambda}A^{\pi}(s,a^{\prime},g))\boldsymbol{d}a^{\prime}.$
@@
$=\arg\min_{\pi}\mathbb{E}_{s,g\sim{\mathcal{B}_{r}}}\mathbb{E}_{a\sim{\mathcal{B}_{r}}}\big{[}-\pi(a|s,g)$
$\cdot\exp(\frac{1}{\lambda}A^{\pi}(s,a,g))\big{]}$ (15)
@@
$\mathcal{J}_{wbc}(\pi)=\mathbb{E}_{s,a,g\sim{\mathcal{B}_{r}}}\left[\exp(\frac{1}{\lambda}A^{\pi}(s,a,g)\cdot\log(\pi(a|s,g)))\right].$ (17)
@@ A.3 Baseline Details
    $\min_{\pi}-\mathbb{E}_{(s_{t},a_{t},g)\thicksim\mathcal{B}_{r}}\big{[}\gamma^{i-t}\exp_{clip}(A(s_{t},a_{t},g))\\ \cdot\epsilon(A(s_{t},a_{t},g))\log\pi(a_{t}\mid s_{t},g)\big{]},$
@@
    $\hat{d}(s_{i+1},g)=-\alpha\log\mathbb{E}_{k}\sim p_{\zeta}^{r}(\cdot|(s_{i+1},g)\cdot\left[e^{-k/(I\alpha)}\right],$ (25)
@@
    where $c(s_{i},g)=\mathbb{1}\left[\phi(c_{i+1})\neq g\right]/I$.
@@
    where $\omega$ is the temperature parameter, and $\pi$ is the policy parameterized by $\theta$.
@@ A.4 Additional Experiments Results
Relabeling Ratio
The multi-goal framework we consider assumes that data is associated with goal labels. We explore the influence of the relabeling ratio on performance, as depicted in Figure 6. The results in Figure 6 indicate that Q-WSL remains robust across different relabeling ratios. Furthermore, Q-WSL consistently outperforms competitive methods such as WGCSL and DDPG+HER at various relabeling ratios.
@@
The Impact of Hyperparameter $\eta$
Since our approach optimizes both supervised learning (SL) and reinforcement learning (RL) concurrently, this section explores the influence of the balancing parameter $\eta$. We evaluate $\eta$ values from the set $\{0.1,0.2,1.0,3.0\}$ and compare the results against competitive HER-based algorithms such as WGCSL and DDPG+HER, as shown in Figure 7. The findings in Figure 7 reveal that Q-WSL consistently delivers superior performance over the other algorithms, regardless of the $\eta$ parameter variation. This demonstrates that our method maintains robustness and is not significantly affected by changes in the $\eta$ parameter.